# clem:todd: A Framework for the Systematic Benchmarking of LLM-Based Task-Oriented Dialogue System Realisations

### v1

@@ Abstract
The emergence of instruction-tuned large language models (LLMs) has advanced the field of dialogue systems, enabling both realistic user simulations and robust multi-turn conversational agents. However, existing research often evaluates these components in isolation—either focusing on a single user simulator or a specific system design—limiting the generalisability of insights across architectures and configurations. In this work, we propose clem:todd (chat-optimized LLMs for task-oriented dialogue systems development)  [^1] https://github.com/kranti-up/clemtoddfw, a flexible framework for systematically evaluating dialogue systems under consistent conditions. clem:todd enables detailed benchmarking across combinations of user simulators and dialogue systems, whether existing models from literature or newly developed ones. It supports plug-and-play integration and ensures uniform datasets, evaluation metrics, and computational constraints. We showcase clem:todd’s flexibility by re-evaluating existing task-oriented dialogue systems within this unified setup and integrating three newly proposed dialogue systems into the same evaluation pipeline. Our results provide actionable insights into how architecture, scale, and prompting strategies affect dialogue performance, offering practical guidance for building efficient and effective conversational AI systems.
@@ 1 Introduction
Task-oriented dialogue systems (Budzianowski et al., 2018; McTear, 2002) help users complete specific goals, such as booking travel or making reservations, through multi-turn natural language interactions. These systems must accurately interpret user intent, manage ambiguity, and respond coherently. With the rise of LLMs, these systems (Hudecek and Dusek, 2023; Xu et al., 2024) have significantly improved in handling task-oriented, goal-driven conversations.
@@
In addition to enhancing dialogue system capabilities, LLMs now play dual roles: as end-to-end dialogue systems (Chung et al., 2023; Dong et al., 2025) and as user simulators (Kojima et al., 2022; Kazi et al., 2024) for training and evaluation. However, evaluation remains a challenge (Casas et al., 2020; Braggaar et al., 2023). Current evaluations often suffer from inconsistent datasets, metrics, and compute settings, making it difficult to compare models or draw conclusions about system design.
@@
To address this gap, we introduce clem:todd (see Figure 1), an evaluation framework that builds on the self-play framework setup for TOD. The user simulator and dialogue system act as the two players in the self-play setup, with a central controller (game master) managing turn-taking, task flow, and interactions with external tools such as databases. This setup supports consistent benchmarking, stress testing, and detailed analysis under shared datasets, metrics, and resource constraints.
Using clem:todd, we explore three research questions: (1) Can LLM-based self-play be adapted to evaluate interactive TOD systems? (2) On the MultiWOZ2.2 benchmark — a widely used dataset for multi-domain TOD — how do user simulators and dialogue systems perform across different models and architectures? (3) Can clem:todd facilitate evaluation beyond MultiWOZ to plug in new instances, either derived from other datasets or generated synthetically, to enhance robustness testing?
@@ 2 Related Work
Our work builds on research in four areas: instruction-tuned LLM-based TOD systems, LLM-based user simulators, LLM-based task evaluation, and self-play frameworks for multi-turn interaction.
@@
While ConvLab (Zhu et al., 2020) and ParlAI [^2] https://parl.ai/ offer infrastructure for developing dialogue agents, and DialEvalMetrics (Yeh et al., 2021) focuses on evaluation, none provide a unified setup for controlled, plug-and-play experimentation across LLM-based user simulators and dialogue systems, evaluated with consistent metrics and compute constraints.
@@
Building on these foundations, we propose clemtod, an evaluation framework that integrates LLM-based user simulators, dialogue systems, and automatic evaluation. It supports flexible combinations of monolithic, modular, and multi-agent architectures and evaluates them under consistent metrics and compute settings. This design enables comprehensive benchmarking and reveals system interaction dynamics in controlled conditions.
@@ 3 Methodology
We build on clembench (Chalamalasetti et al., 2023), a framework that realises conversational games between two (LLM-simulated) players coordinated by a Game Master. clem:todd reuses and modifies the two-player setup as: (a) Player A is an LLM-based user simulator that receives a goal and a prompt (see Appendix 7) describing the task; (b) Player B is a dialogue system that receives user simulator utterances and a prompt (see Appendix 8 and 9) describing the task, with the objective of satisfying Player A’s goal; and (c) the Game Master orchestrates the interaction by passing utterances between the two players and managing turn-level coordination.
The Game Master enforces stricter format constraints than the base clembench setup by requiring adherence to a predefined Tool Schema (in JSON format) (e.g., followup, querydb, validatebooking) to reflect the demands of real-world deployment. Any violation aborts the conversation, reinforcing that instruction following and format compliance is essential. Valid tool calls are executed programmatically, either to retrieve database results, validate booking details, or forward follow-up messages to the user simulator. The schema is given in Figures 14-19.
@@ 3.2 User Simulator
We leverage the generative capabilities of LLMs to simulate realistic user interactions. The user simulator is prompted (see Figure 7 in Appendix A.1) with a set of goals, represented as natural language objectives, and engages in a dialogue with the system under evaluation to fulfill them. The prompt includes relevant task context and dialogue history, enabling the simulator to produce coherent, goal-oriented responses throughout the conversation.
@@ 3.3.1 Existing Dialogue Systems
We integrated two representative systems from prior work: AutoTOD, a zero-shot monolithic system that reported results on MultiWOZ 2.0 dataset (Xu et al., 2024), and a modular pipeline (Hudecek and Dusek, 2023) that reported results on MultiWOZ 2.2 dataset. These systems were selected for their architectural diversity and publicly available implementations. Attempts to incorporate other approaches weren’t successful  [^3] InstructTOD, evaluted on Multiwoz 2.1 (Chung et al., 2023) encountered compatibility issues between the agent implementation and LLM outputs, which led to repeated execution errors. For ProTOD, evaluted on Multiwoz 2.0 (Dong et al., 2025), the publicly released code lacked essential implementation components, rendering integration into our framework infeasible. (see Appendix A.2 for details).
@@ 3.3.2 Proposed Dialogue Systems
This design decomposes the dialogue pipeline processing into sub-modules for intent detection, slot extraction, dialogue management, and response generation (see Figures 11, 12, 13 for prompt templates). We explore two variants that reflect distinct control strategies: (a) Modular-Prog, where a programmatic dialogue manager executes modules in a fixed order (a static pipeline); and (b) Modular-LLM, where an LLM acts as the dialogue manager (see Figure 9 for the prompt), dynamically selecting the next sub-module based on intermediate outputs (an adaptive pipeline). All sub-modules use LLMs and produce outputs aligned with the Tool Schema. Outputs are validated before being passed downstream, and dialogue context is maintained across turns. Both variants follow a simple, traditional modular design, allowing us to assess how classical pipelines perform when augmented with LLMs.
@@ 4 Experimental Setup
Model Naturalness Coherence Dialogue Diversity
Llama-3.2-1B 3.84 2.84 1.00
Llama-3.2-3B 2.42 1.75 1.00
Qwen2.5-7B 2.75 2.42 1.05
Llama-3.1-8B 2.82 2.18 1.02
Qwen2.5-32B 4.65 2.95 1.00
Llama-3.3-70B 4.42 2.97 1.00
Table 1: Dialogue quality comparison of user simulators for the Monolithic architecture-based dialogue system (using the model: Qwen2.5-32B), evaluated on Naturalness (N), Coherence (C), and Dialogue Diversity (D) metrics using GPT-4o. Higher scores indicate better performance.
@@ 5.1.1 Analyzing the Role of User Simulators
Table 2: Task success (booking) accuracy across different LLM pairings as user simulators (US) and dialogue systems (DS) in the clemtod framework. The rows show the models used as US, while the columns are the models taking the DS side, in the variants Monolithic (M), Modular-Programmatic (MP), and Modular-LLM (ML) architecture.
To identify user simulators that shows coherent and natural behavior, we evaluate dialogue systems of varying capacities ($32$B and $70$B) paired with simulators from the Qwen and LLaMA model families, ($1$B $\sim$ $70$B) on single-domain test dialogues.
@@
Table 3: Performance of various dialogue system architectures within the clemtod framework. The table presents inform(I), and booking accuracy (B) for Monolithic, Modular-Prog, and Modular-LLM architectures across different LLMs. Systems based on Xu et al. and H et al. are also evaluated within the clemtod setup. Larger models consistently yield better performance.Hudecek and Dusek (2023)’s code do not support task success computation.
We assess dialogue quality using both automatic and human evaluations. For automatic evaluation, we adopt the LLM-as-a-judge framework (Kazi et al., 2024), using GPT-4o in a zero-shot setting to rate generated user utterances. Results in Table 1 indicate that both naturalness and coherence generally improve with increasing model size, while dialogue-level diversity remains relatively constant $\sim 1.0$). These trends are corroborated by parallel evaluations using the open-weight Llama-3.3-70B model as the judge (see Table 11).
To further validate these findings, we conducted a “Turing Test” (see Appendix A.4) where an annotator choose the more natural dialogue between an LLM output and its corresponding ground-truth. We randomly selected 50 dialogues. $19$ out of $50$ sampled dialogues generated by Qwen2.5-32B were preferred over the corpus version, compared to $8$ from LLaMA-3.3-70B. This shows Qwen2.5-32B’s strength in generating realistic user behaviour.
Although dialogue quality assessment provides insight into simulated dialogues, it is also important to evaluate the robustness of dialogue systems to user simulators for reliable evaluation. As shown in Table 2, the overall performance of dialogue systems generally is sensitive to the choice (and with this, as established by Table 1, the quality) of the user simulator, but to differing extents: The dialogue systems realised with Qwen2.5-32B appear to be more capable of maintaining task success despite differences in simulator behavior. These systems can recover from incoherent user turns and use clarifying strategies to manage imperfect inputs and maintain goal alignment. In contrast, task success declines for LLaMA-3.3-70B when the user simulator is too small or behaves inconsistently.
[…]
5.1.2 Evaluating Dialogue Systems
@@
Table 4: Cost comparison of dialogue system architectures within the clemtod framework. The table reports token-based cost (T) and FLOPS cost (F), both measured in USD (\$) per dialogue, for Monolithic, Modular-Prog, and Modular-LLM architectures across different LLMs.
In contrast, open-weight models perform better in monolithic configurations, achieving booking success rates around $\sim 0.77$, but show significantly reduced performance in modular setups. Most failures stem from invalid response formats or prematurely signalling the end of conversation without conversing about all the goals (see Appendix 5). In modular configurations, the overall task is decomposed into subtasks, and individual modules operate without access to the full dialogue history. This limited context appears to hinder the model’s ability to consistently satisfy goal constraints, reducing task success. In monolithic setups, a detailed analysis shows that these models can reach booking rates as high as $0.95$ (Qwen2.5–32B) for single-domain dialogues (see Table 13 for the breakdown of results across domains). However, their overall performance declines in multi-domain scenarios, where increased complexity and the need for cross-domain goal tracking present greater challenges.
@@ 5.2 Evaluating Computational Efficiency of Dialogue Systems
Figure 3: Cost breakdown of dialogue system architectures across models in the clemtod framework. The plot shows token cost (dashed lines, left y-axis) and FLOPs cost (solid lines with markers, right y-axis) per dialogue for Monolithic, Modular-Prog, and Modular-LLM systems.
As shown in Table 4 and visualized in Figure 3, the Monolithic architecture has the lowest computational cost due to its single-pass execution, avoiding repeated LLM calls. In contrast, Modular Prog and Modular LLM incur higher costs. Modular Prog has the highest FLOP cost, up to \$$\$1.33$ with Llama-3.3-70B, $\sim 2$ times more than its Monolithic counterpart, reflecting poorly on efficiency.
@@ 5.2.1 Performance-Cost Trade-offs
Overall, architecture selection should align with application-specific priorities. Monolithic setups are preferable when minimizing compute or API usage is critical. Modular architectures—particularly Modular-LLM—offer a compelling compromise.

### v2

@@ Abstract
The emergence of instruction-tuned large language models (LLMs) has advanced the field of dialogue systems, enabling both realistic user simulations and robust multi-turn conversational agents. However, existing research often evaluates these components in isolation, either focusing on a single user simulator or a specific system design, limiting the generalisability of insights across architectures and configurations. In this work, we propose clem:todd (chat-optimized LLMs for task-oriented dialogue systems development), a flexible framework for systematically evaluating dialogue systems under consistent conditions. clem:todd enables detailed benchmarking across combinations of user simulators and dialogue systems, whether existing models from literature or newly developed ones. To the best of our knowledge, clem:todd is the first evaluation framework for task-oriented dialogue systems that supports plug-and-play integration and ensures uniform datasets, evaluation metrics, and computational constraints. We showcase clem:todd’s flexibility by re-evaluating existing task-oriented dialogue systems within this unified setup and integrating three newly proposed dialogue systems into the same evaluation pipeline. Our results provide actionable insights into how architecture, scale, and prompting strategies affect dialogue performance, offering practical guidance for building efficient and effective conversational AI systems.
@@ 1 Introduction
Task-oriented dialogue systems (McTear, 2002; Li et al., 2017; Eric et al., 2017; Budzianowski et al., 2018; Balakrishnan et al., 2019; Chen et al., 2019; Elder et al., 2020) help users complete specific goals, such as booking travel or making reservations, through multi-turn natural language interactions. These systems must accurately interpret user intent, manage ambiguity, and respond coherently.
@@
With the rise of LLMs, these systems (Hudecek and Dusek, 2023; Xu et al., 2024) have significantly improved in handling task-oriented, goal-driven conversations. In addition to enhancing dialogue system capabilities, LLMs now play dual roles: as end-to-end dialogue systems (Chung et al., 2023; Dong et al., 2025) and as user simulators (Kojima et al., 2022; Kazi et al., 2024) for training and evaluation. However, evaluation remains a challenge (Casas et al., 2020; Braggaar et al., 2023). Current evaluations often suffer from inconsistent datasets, metrics, and compute settings, making it difficult to compare models or draw conclusions about system design.
@@
To address this gap, we introduce clem:todd  [^1] https://github.com/clp-research/clem-todd (see Figure 1), an evaluation framework that builds on the self-play framework setup for TOD. The user simulator and dialogue system act as the two players in the self-play setup, with a central controller (game master) managing turn-taking, task flow, and interactions with external tools such as databases. This setup supports consistent benchmarking, stress testing, and detailed analysis under shared datasets, metrics, and resource constraints.
Using clem:todd, we explore three research questions: (1) Can LLM-based self-play be adapted to evaluate interactive TOD systems? (2) On the MultiWOZ2.2 benchmark, a widely used dataset for multi-domain TOD, how do user simulators and dialogue systems perform across different models and architectures? (3) Can clem:todd facilitate evaluation beyond MultiWOZ to plug in new instances, either derived from other datasets or generated synthetically, to enhance robustness testing?
@@ 2 Related Work
Our work builds on research in: LLM-based TOD systems, user simulators, task evaluation, and self-play frameworks for multi-turn interaction.
@@
Ghandeharioun et al. (2019) propose an evaluation framework for dialogue-system self-play in open-domain dialogue systems. Cheng et al. (2022) discuss the importance of using a goal-oriented user simulator for improved evaluation. Xu et al. (2024) showcase the evaluation of different dialogue systems against a single user simulator. Additionally ConvLab (Zhu et al., 2020) and ParlAI [^2] https://parl.ai/ offer infrastructure for developing dialogue agents, while DialEvalMetrics (Yeh et al., 2021) focuses on evaluation. However, none provide a unified setup for controlled, plug-and-play experimentation across LLM-based user simulators and dialogue systems, evaluated with consistent metrics and compute constraints. clem:todd fills this gap by enabling systematic exploration of architectural choices, simulator–system interactions, and scaling trade-offs.
@@
Building on these foundations, we propose clem:todd  an evaluation framework that integrates LLM-based user simulators, dialogue systems, and automatic evaluation. It supports flexible combinations of monolithic, modular, and multi-agent architectures and evaluates them under consistent metrics and compute settings. This design enables comprehensive benchmarking and reveals system interaction dynamics in controlled conditions.
@@ 3 Methodology
We build on clembench (Chalamalasetti et al., 2023), a framework that realises conversational games between two (LLM-simulated) players coordinated by a Game Master. clem:todd reuses and modifies the two-player setup as: (a) Player A is an LLM-based user simulator that receives a goal and a prompt (see Appendix 6) describing the task; (b) Player B is a dialogue system that receives user simulator utterances and a prompt (see Appendix 7 and 8) describing the task, with the objective of satisfying Player A’s goal; and (c) the Game Master orchestrates the interaction by passing utterances between the two players and managing turn-level coordination.
The Game Master enforces stricter format constraints than the base clembench setup by requiring adherence to a predefined Tool Schema (in JSON format) (e.g., followup, querydb, validatebooking) to reflect the demands of real-world deployment. Any violation aborts the conversation, reinforcing that instruction following and format compliance is essential. Valid tool calls are executed programmatically, either to retrieve database results, validate booking details, or forward follow-up messages to the user simulator. The schema is given in Figures 13-18 in Appendix.
@@
clem:todd draws inspiration from “self-play” and adapts it to goal-driven task-oriented dialogue systems, where the user and system take on distinct roles. Building on prior works (Zhu et al., 2020; Xu et al., 2024; Kazi et al., 2024), clem:todd employs a schema-constrained interaction loop and supports plug-and-play integration of user simulators and dialogue systems across different architectures, enabling more robust and extensible benchmarking.
@@ 3.2 User Simulator
We leverage the generative capabilities of LLMs to simulate realistic user interactions. The user simulator is prompted (see Figure 6 in Appendix A.1) with a set of goals, represented as natural language objectives, and is intended to engage in a dialogue with the system under evaluation to fulfill them. The prompt includes relevant task context and dialogue history, enabling the user simulator to produce coherent, goal-oriented responses in the conversation. Such simulators facilitate automated evaluation workflows, complementing real users.
@@ 3.3.1 Existing Dialogue Systems
We integrated two representative systems from prior work: AutoTOD, a zero-shot monolithic system that reported results on MultiWOZ 2.0 dataset (Xu et al., 2024), and a modular pipeline (Hudecek and Dusek, 2023) that reported results on MultiWOZ 2.2 dataset. These systems were selected for their architectural diversity and publicly available implementations. Attempts to incorporate other approaches were not successful 333 InstructTOD, evaluted on Multiwoz 2.1 (Chung et al., 2023) encountered compatibility issues between the agent implementation and LLM outputs, which led to repeated execution errors. For ProTOD, evaluted on Multiwoz 2.0 (Dong et al., 2025), the publicly released code lacked essential implementation components, rendering integration into our framework infeasible. (see Appendix A.2 for details).
@@ 3.3.2 Proposed Dialogue Systems
This design decomposes the dialogue pipeline processing into sub-modules for intent detection, slot extraction, dialogue management, and response generation. (see Figures 10, 11, 12 in the Appendix for prompt templates). We explore two variants that reflect distinct control strategies: (a) Modular-Prog, where a programmatic dialogue manager executes modules in a fixed order (a static pipeline); and (b) Modular-LLM, where an LLM acts as the dialogue manager, (see Figure 8 for the prompt), dynamically selecting the next sub-module based on intermediate outputs (an adaptive pipeline). All sub-modules use LLMs and produce outputs aligned with the Tool Schema. Outputs are validated before being passed downstream. Both variants follow a simple, traditional modular design, allowing us to assess how classical pipelines perform when augmented with LLMs.
@@ 5.1 Extending Self-Play to Task-Oriented Dialogue
Model (US) Naturalness Coherence Diversity TT
Llama-3.2-1B 3.84 2.84 1.00 -
Llama-3.2-3B 2.42 1.75 1.00 -
Qwen2.5-7B 2.75 2.42 1.05 -
Llama-3.1-8B 2.82 2.18 1.02 -
Qwen2.5-32B 4.65 2.95 1.00 0.38
Llama-3.3-70B 4.42 2.97 1.00 0.16
Table 1: Dialogue quality comparison of user simulators for the Monolithic architecture-based dialogue system (using the model: Qwen2.5-32B), evaluated on Naturalness (N), Coherence (C), and Dialogue Diversity (D) metrics using GPT-4o. The Turing Test (TT) metric measures the percentage of dialogues from a random sample of $50$ that are judged as human-like, based on human evaluation. Higher scores indicate better performance.
@@ 5.1.1 Analyzing the Role of User Simulators
To identify user simulators that show coherent and natural behavior, we evaluate dialogue systems of varying capacities ($32$B and $70$B) paired with simulators from the Qwen and LLaMA model families, ($1$B $\sim$ $70$B) on single-domain test dialogues.
We assess dialogue quality using both automatic and human evaluations. For automatic evaluation, we adopt the LLM-as-a-judge framework (Kazi et al., 2024), using GPT-4o in a zero-shot setting to rate generated user utterances. Results in Table 1 indicate that both naturalness and coherence generally improve with increasing model size, while dialogue-level diversity remains relatively constant ( $\sim 1.0$). These trends are corroborated by parallel evaluations using the open-weight Llama-3.3-70B model as the judge (see Table 12).
To further validate these findings, we conducted a “Turing Test” (see Appendix A.4) where an annotator chose the more natural dialogue between an LLM output and its corresponding ground-truth. This evaluation assesses the dialogues produced by Qwen2.5-32B and LLaMA-3.3-70B relative to ground-truth dialogues for the same task, focusing specifically on perceived naturalness.
We developed a simple user interface (see Figure 5 in Appendix) that displays two dialogues side by side, allowing annotators to select which dialogue appears more natural. In each comparison, one dialogue is generated by either Qwen2.5-32B or LLaMA-3.3-70B, and the other is the corresponding ground-truth dialogue from the corpus (Budzianowski et al., 2018).
As shown in Table 1, $19$ out of $50$ randomly sampled dialogues generated by Qwen2.5-32B were preferred ($0.38$) over the ground truth, compared to $8$ ($0.16$) from LLaMA-3.3-70B. This result highlights Qwen2.5-32B’s strength in generating realistic user behavior.
5.1.2 Robustness of Dialogue Systems to User Simulator Variability
How robust are dialogue systems when interacting with user simulators of varying capability and coherence? Although dialogue quality assessment (in Section 5.1.1) provides insight into simulated dialogues, it is also important to evaluate the robustness of dialogue systems to user simulators for reliable evaluation. As shown in Table 2, the overall performance of dialogue systems generally is sensitive to the choice (and with this, as established by Table 1, the quality) of the user simulator, but to differing extents: The dialogue systems realised with Qwen2.5-32B appear to be more capable of maintaining task success despite differences in simulator behavior. These systems can recover from incoherent user turns and use clarifying strategies to manage imperfect inputs and maintain goal alignment. In contrast, task success declines for LLaMA-3.3-70B when the user simulator is too small or behaves inconsistently.
[…]
@@ 5.1.2 Robustness of Dialogue Systems to User Simulator Variability
User Spread 0.58 0.46 0.61 0.62 0.25 0.62
Table 2: Task success (booking) accuracy across different LLM pairings as user simulators (US) and dialogue systems (DS) in the clem:todd framework. Rows indicate the models used as US, and columns show the DS models evaluated in three architectural variants: Monolithic (M), Modular-Programmatic (MP), and Modular-LLM (ML). The bottom-most row (User Spread) reports the standard deviation across user simulators for each DS configuration, reflecting its sensitivity to variation in simulator behavior. Lower User Spread values indicate greater robustness, but should be considered alongside overall task success.
5.1.3 Evaluating Dialogue Systems
@@
Table 3: Performance of various dialogue system architectures within the clem:todd framework. The table presents inform(I), and booking accuracy (B) for Monolithic, Modular-Prog, and Modular-LLM architectures across different LLMs. Systems based on Xu et al. (2024) and Hudecek and Dusek (2023) are also evaluated within the clem:todd setup. Larger models consistently yield better performance. Hudecek and Dusek (2023)’s system does not support booking accuracy computation.
@@
In contrast, open-weight models perform better in monolithic configurations, achieving booking success rates around $\sim 0.77$, but show significantly reduced performance in modular setups. Most failures stem from invalid response formats or prematurely signalling the end of conversation without conversing about all the goals (see Appendix 4). In modular configurations, the overall task is decomposed into subtasks, and individual modules operate without access to the full dialogue history. This limited context appears to hinder the model’s ability to consistently satisfy goal constraints, reducing task success. In monolithic setups, a detailed analysis shows that these models can reach booking rates as high as $0.95$ (Qwen2.5–32B) for single-domain dialogues (see Table 13 for the breakdown of results across domains). However, their performance drops in multi-domain scenarios due to increased complexity and cross-domain goal tracking.
@@
Table 4: Cost comparison of dialogue system architectures within the clem:todd framework. The table reports token-based cost (T) and FLOPS cost (F), both measured in USD (\$) per dialogue, for Monolithic, Modular-Prog, and Modular-LLM architectures across different LLMs.
@@ 5.2 Evaluating Computational Efficiency of Dialogue Systems
Model Monolithic Modular-Prog Modular-LLM
Llama-3.2-1B 18.85 1.61 18.72
Llama-3.2-3B 5.21 8.32 16.56
Qwen2.5-7B 6.69 11.38 13.70
Llama-3.1-8B 5.97 11.16 42.08
Qwen2.5-32B 8.52 10.16 7.17
Llama-3.3-70B 25.33 34.07 47.98
GPT-4o 4.10 8.61 4.37
Table 5: Latency time (in seconds) comparison across dialogue system architectures in the clem:todd framework. The table reports end-to-end response time for Monolithic, Modular-Prog, and Modular-LLM architectures evaluated across different LLMs.
@@
As shown in Table 4, the Monolithic architecture has the lowest computational cost due to its single-pass execution, avoiding repeated LLM calls. In contrast, Modular Prog and Modular LLM incur higher costs. Modular Prog has the highest FLOP cost, up to \$$\$1.33$ with Llama-3.3-70B, $\sim 2$ times more than its Monolithic counterpart, reflecting poorly on efficiency.
Additionally, we computed the latency time for each system, defined as the total time elapsed from the first interaction initiated by the user simulator until the task completion or the maximum number of dialogue turns ($15$) is reached. As shown in Table 5, Monolithic systems exhibit the lowest latency times across most configurations. This is expected, as Monolithic systems involve a single model call per turn, whereas the Modular-Prog and Modular-LLM systems require multiple calls to sub-modules, increasing overall latency. The highest latency is observed for Llama-3.3-70B, which was executed on two GPUs, suggesting that model size and hardware configuration contribute significantly to response time. Among the modular variants, Modular-LLM systems generally exhibit higher latency than Modular-Prog systems, which can be attributed to the overhead of using an LLM as the dialogue manager in addition to the other modules. In contrast, the lowest latency is observed for Llama-3.2-1B, which is primarily due to early termination of many dialogues resulting from format violations leading to fewer turns and thus shorter overall latency.
@@
Overall, architecture selection should align with application-specific priorities. Monolithic setups are preferable when minimizing compute or API usage is critical. Modular architectures, particularly Modular-LLM, offer a compelling compromise.